### v1

# Adapting Multilingual Embedding Models to Historical Luxembourgish

Abstract
The growing volume of digitized historical texts requires effective semantic search using text embeddings. However, pre-trained multilingual models, typically evaluated on contemporary texts, face challenges with historical digitized content due to OCR noise and outdated spellings. We explore the use of multilingual embeddings for cross-lingual semantic search on historical Luxembourgish, a low-resource language. We collect historical Luxembourgish news articles spanning various time periods and use GPT-4o to segment and translate them into closely related languages, creating 20,000 parallel training sentences per language pair. We further create a historical bitext mining evaluation set and find that these models struggle to perform cross-lingual search on historical Luxembourgish. To address this, we propose a simple adaptation method using in-domain training data, achieving up to 98% accuracy in cross-lingual evaluations. We release our adapted models and historical Luxembourgish-German/French bitexts to support further research.
Adapting Multilingual Embedding Models to Historical Luxembourgish
Andrianos Michail University of Zurich andrianos.michail@cl.uzh.ch Corina Julia Raclé University of Zurich corinajulia.racle@uzh.ch
Juri Opitz University of Zurich jurialexander.opitz@cl.uzh.ch Simon Clematide University of Zurich simon.clematide@cl.uzh.ch
1 Introduction
The ability to access and explore historical texts such as newspapers has advanced rapidly through the digitization efforts of libraries and archives. This exploration has traditionally been accomplished through keyword-based searches, enhanced by semantic enrichment such as named entities.
More recently, vector-based semantic search has gained prominence through embedding models, which are also widely used in Retrieval-Augmented Generation to increase response factuality. In embedding benchmarks (Muennighoff et al., 2023), massively multilingual embedding models such as Multilingual E5 and GTE Multilingual (Zhang et al., 2024; Wang et al., 2024), trained on diverse multilingual corpora, have demonstrated strong performance in multilingual and cross-lingual semantic search across a wide range of languages. However, for low-resource languages such as Luxembourgish (LB), where multilingual models have limited exposure, their performance remains uncertain. When applied to semantic search in imperfectly digitized historical collections, these models must be robust to OCR errors and old spellings. The mismatch between this domain and the clean, contemporary texts on which these multilingual models were trained, along with their already limited support for Luxembourgish, poses a challenge for developing effective exploration tools for historical Luxembourgish newspaper archives.
To address this issue, we compile 2,340 historical Luxembourgish news articles from different time periods and use GPT-4o to segment and translate them into contemporary German (DE) and French (FR). The resulting parallel sentences serve as fine-tuning data to adapt existing multilingual embedding models for imperfectly digitized historical Luxembourgish.
Our contributions:
1. To adapt multilingual embeddings for digitized historical Luxembourgish, we propose to build training data through a prompt-based translation approach with GPT-4o.
2. For evaluation, we frame a historical Bitext Mining tasks and create a high-quality test set containing 233 source news articles (LB-DE: 2,127 sentences; LB-FR: 2,157 sentences).
3. To address this task and assess our proposed adaptation method, we evaluate models based on LaBSE (Feng et al., 2022), M-GTE (Zhang et al., 2024) and LuxEmbedder (Philippy et al., 2025), and analyze potential trade-offs in performance.
We will release our code and datasets publicly.
2 Related Work
The multilingual bitext mining model LaBSE (Feng et al., 2022) is trained with translation ranking loss and negative samples. It has been trained roughly on less than 100 Luxembourgish-English sentence pairs and specializes in zero-shot bitext mining. A recent model is GTE Multilingual (M-GTE) (Zhang et al., 2024), a multilingual embedding model designed for long context text representation and reranking. M-GTE has been trained with hard negatives and has included 50,000 Luxembourgish pairs within its contrastive pre-training.
Specific model adaptations to Luxembourgish have also been developed. One example is LuxemBERT (Lothritz et al., 2022), a monolingual BERT model pre-trained for Luxembourgish using augmented data, partially generated by translating texts from closely related languages and incorporating relevant text sources.
Closely related to our work, LuxEmbedder (Philippy et al., 2025) used OpenAI’s text-embedding-3-small and LaBSE to bitext mine a set of 20,000 parallel sentences for each language pair between Luxembourgish, English, and French. These parallel sentences were then used to further fine-tune LaBSE, improving performance on contemporary Luxembourgish evaluation sets. However, its ability to handle Luxembourgish texts from different historical periods—potentially affected by digitization errors common in large-scale historical text collections—remains unclear.
3 Method
For training and evaluating embedding models for digitized historical Luxembourgish news articles, we create parallel texts.
3.1 Parallel Historical Luxembourgish
Model Historical Bitext Mining Contemporary LB Domain Adapted LBDE LBFR SIB 200 (LB) PARALUX OpenAI/text-embedding-3-small – 82.33 78.36 40.20 15.71 OpenAI/text-embedding-3-large – 88.15 86.18 58.82 26.28 LaBSE – 94.01 93.34 43.24 38.14 LuxEmbedder – 85.48 85.07 65.59 52.24 M-GTE – 87.55 85.71 55.78 70.51 LaBSE (ours) LBDE 98.52 97.98 34.02 14.70 LBFR 98.24 97.88 35.88 25.96 LBDE/FR 98.73 98.21 36.27 20.19 LuxEmbedder (ours) LBDE 98.38 97.72 50.10 25.96 LBFR 97.89 97.74 55.88 25.32 LBDE/FR 98.52 97.91 43.04 24.68 M-GTE (ours) LBDE 97.34 96.25 44.41 41.99 LBFR 96.26 95.78 59.71 50.32 LBDE/FR 97.37 96.30 49.22 44.55 Random Baseline – 0.00 0.00 14.28 50.00
We build our translated parallel data sets LB-DE and LB-FR from monolingual Luxembourgish texts sourced from the publicly available BNL newspaper archive.111https://data.bnl.lu/data/historical-newspapers/ Our data consists of articles from newspapers published between 1841 and 1948. To select diverse samples for translation, we first cluster the articles into 2,000 groups by K-Mean and keep the 605 clusters with more than 20 articles. We select articles through a two-step process, resulting in a total of 2,340 articles, as shown in Figure 1. First, we retrieve the most representative article from each cluster, ensuring it contains between 5 and 20 sentences. In a second round, we randomly sample three additional articles per cluster under the same length conditions.
We prompt GPT-4o to segment historical Luxembourgish articles and generate sentence-level translation pairs separately for German and French (see Prompt 3). The model is instructed to preserve the original meaning and structure as closely as possible while reconstructing sentences affected by OCR errors that could hinder translation. This process yields approximately 22,500 sentence pairs for both LB-DE and LB-FR. Notably, the model’s consistent sentence segmentation results in 74.9% of sentences forming triplets across the three languages.
To ensure fidelity to the original articles, we calculate the percentage of regenerated Luxembourgish sentences that do not exactly match their source texts. These account for 1.5% of all historical Luxembourgish sentences per language pair, which we manually correct. Most mismatches result from missing or added punctuation, modernized spelling, and, in rare cases, errors caused by the LLM not adhering to the instructed format.
To assess translation quality, a trilingual native LB speaker annotates 100 randomly sampled sentence triplets after removing 15 samples with obvious OCR issues or other factors. Of 100 Luxembourgish sentences (presented without context), 86 are rated comprehensible or at least confidently guessable (24). 14 are incomprehensible in isolation due to OCR errors and old spellings. The translations of the incomprehensible sentences are not assessed. The German translations of the comprehensible and confidently guessable sentences are rated as adequate in 77 cases (89.5%): having missing details in 8 cases and only 1 case of insufficient translation. For the French translations, 76 cases are adequate (88.4%), 9 are missing minor details, and 1 is of insufficient quality. Samples from our dataset including their annotation labels are available in the Appendix (Table 2).
3.2 Framing an Evaluation Task: Historical Bitext Mining
From our parallel dataset, we reserve a held-out test set of 233 articles (2,127 sentences) to construct a historical LB-to-DE/FR Bitext Mining evaluation set. The evaluation method requires the model to assign higher similarity to the correct parallel sentence than to any of the 2,126 other candidates in the set in order to count as a true positive. We report bi-directionally averaged accuracy. To minimize false negatives stemming from near-identical sentences, we exclude candidate sentences with a Levenshtein similarity score above 0.85 to the source sentence after removing non-alphanumeric characters from both. This filtering affects 56 source-candidate pairs (2.6% of the source sentences) in German and 64 source-candidate pairs (2.97% of the sentences) in French. A human check of the excluded pairs at different thresholds confirms the appropriateness of this filtering process and the chosen threshold.
3.3 Contemporary LB Evaluation Datasets
To contextualize the evaluation on our targeted in-domain data, we replicate two evaluation methods of Philippy et al. (2025) for contemporary Luxembourgish. First, PARALUX is a monolingual Luxembourgish paraphrase detection test set designed to assess embedding models. Performance is measured by how often an embedding model assigns higher similarity to an anchor-positive pair than to an anchor-negative pair. The negative sentences are adversarially generated to ensure high lexical similarity, and human-corrected to ensure that they are true negatives. Second, SIB 200 (LB) is a repurposed dataset sourced from the ‘Flores’-dataset (Costa-jussà et al., 2022), framed as a monolingual zero-shot classification task based on embedding similarity. For further details, we refer to Philippy et al. (2025).
3.4 Adapting Multilingual Embedding Models to Historical Luxembourgish
We take the aforementioned LaBSE, LuxEmbedder and M-GTE and adapt them to historical Luxembourgish. We fine-tune these embedding models using a MultipleNegativesRankingLoss (Henderson et al., 2018) with batch size 8 over 1 epoch. We experiment with three fine-tuning data variants: (1) LBDE with 20,000 historical Luxembourgish sentences paired with German, (2) LBFR pairs, and (3) LBDE/FR, using 10,000 sentences from each language in mixed batches.
4 Results
We show the results of various models in Table 1.
Historical Bitext Mining – Cross-Lingual Semantic Search
Among off-the-shelf models, LaBSE is the strongest model in both languages. Surprisingly, LuxEmbedder, the Luxembourgish-tuned version of LaBSE, shows an 8.5pp performance drop for both language pairs in our bitext mining task, despite improved performance on contemporary Luxembourgish tasks. Similarly, M-GTE underperforms LaBSE by 5.5pp. Both OpenAI embedding models demonstrate moderate performance in all tasks.
For domain-adapted models, performance improves significantly in historical bitext mining for both language pairs after training on 20,000 parallel sentences, whether using LBDE or LBFR data. However, the best results are consistently achieved with the 20,000 LBDE/FR sentence pairs, yielding higher performance across all models. Notably, after domain adaptation, LuxEmbeddder nearly matches the adapted LaBSE, achieving over 98% accuracy and closing the performance gap observed in the off-the-shelf models. In contrast, M-GTE, the general embedding model, appears to plateau at lower performance levels, trailing the specialized domain-adapted bitext mining model by 1.5pp.
Contemporary vs. historic Luxembourgish: A notable trade-off
While our domain-adapted models improve the embedding of historical Luxembourgish for cross-lingual semantic search in contemporary German and French (as shown on the right side of the Table 1), their performance on monolingual contemporary tasks, particularly PARALUX, declines noticeably after domain adaptation. Interestingly, this degradation is more pronounced when models are trained on the linguistically closer LBDE pairs than LBFR pairs.
The learning curves in Figure 2 illustrate the continuous decline in monolingual performance. Notably, most cross-lingual performance gains occur within the first 2,500 samples, before monolingual Luxembourgish performance declines more sharply. This highlights the trade-off between optimizing for historical cross-lingual tasks and maintaining monolingual performance, as seen in the weaker monolingual results of the best off-the-shelf LaBSE model.
5 Conclusions
In this work, we present a case study on embedding imperfectly digitized historical Luxembourgish texts. As off-the-shelf models prove suboptimal for this task, we propose a method to synthesize reliable data for adapting embedding models to this domain. We generate parallel sentence-segmented documents by translating historical Luxembourgish newspaper articles into two related higher-resource languages (German and French) using GPT-4o as a machine translation system.
To establish a hold-out test set, we reserve 233 articles and frame a bitext mining task for retrieving historical Luxembourgish sentences. Evaluating various multilingual embedding models on this task, we demonstrate that our adaptation approach improves performance by up to 13 percentage points. While our adapted models perform effectively across different historical periods, this improvement comes at the cost of reduced performance in natural language understanding (NLU) for clean contemporary texts. This trade-off highlights the need for future research into embedding historical digitized texts affected by significant OCR errors.
Limitations
Our findings present both positive and negative aspects. On the one hand, our method demonstrates clear benefits for the targeted use case, effectively embedding heterogeneous digitized historical texts and revealing shortcomings in off-the-shelf models. On the other hand, performance drops in control tasks on contemporary texts raise concerns. Does improved robustness to digitization errors come at the cost of weaker language understanding and reduced sensitivity to semantic nuances? Can future training strategies mitigate this trade-off? These important questions remain open for future research.
Acknowledgements
This research is conducted under the project Impresso – Media Monitoring of the Past II Beyond Borders: Connecting Historical Newspapers and Radio. Impresso is a research project funded by the Swiss National Science Foundation (SNSF 213585) and the Luxembourg National Research Fund (17498891).
Appendix A Appendix
| Newspaper | Year | Sentence | LB Comprehensibility Rating | DE Translation Rating | FR Translation Rating |
| D’Union | 1946 |
LB: Si bestét aus 18 000 tuben, weit 30 tonnen a kascht 400 000 dollar.
DE: Sie besteht aus 18.000 Röhren, wiegt 30 Tonnen und kostet 400.000 Dollar. FR: Elle est composée de 18 000 tubes, pèse 30 tonnes et coûte 400 000 dollars. |
Comprehensible | Adequate | Adequate |
| Ons Jorgen | 1946 |
LB: „Daf net, mais ech hu kc properen Teller me’,"
DE: „Das nicht, aber ich habe keinen sauberen Teller mehr. FR: "Certainement pas, mais je n’ai plus d’assiette propre," |
Comprehensible | Adequate | Adequate |
| De Letzeburger | 1893 |
LB: De Batti: Elo hätte mer d’Stemmung gut eriwer, hätte mer elo och nach Rén.
DE: Der Batti: Jetzt hätten wir die Stimmung gut geschafft, hätten wir jetzt auch noch Regen. FR: Le Batti : Maintenant, nous aurions bien passé l’ambiance, si seulement nous avions aussi encore Rén. |
Confidently Guessable | Adequate | Adequate |
| De Letzeburger | 1905 |
LB: Op d’Weis: Das ist im Lehen hfisslich eingerichtet.
DE: Zur Melodie: Das ist im Leben hässlich eingerichtet. FR: À la manière de : Cela est arrangé vilaine dans la vie. |
Confidently Guessable | Adequate | Adequate |
| De Letzeburger | 1893 |
LB: Wann och d’Liss’ché wéss ze feischtren.
DE: Wenn auch die Liss’ché weiß zu feilschen. FR: Même si Liss’ché sait lutter. |
Incomprehensible | / | / |
| Obermosel-Zeitung | 1924 |
LB: In vielen vorkern Bincl alle Krank, 80 cla,BB «lie ?eläer nicdt deBtellt terrien Können.
DE: In vielen Dörfern sind alle krank, so dass die Felder nicht bestellt werden können. FR: Dans de nombreux villages, tous sont malades, si bien que les champs ne peuvent pas être cultivés. |
Incomprehensible | / | / |

### v3

@@
 The growing volume of digitized historical texts requires effective semantic search using text embeddings.
-However, pre-trained multilingual models, typically evaluated on contemporary texts, face challenges with historical digitized content due to OCR noise and outdated spellings.
-We explore the use of multilingual embeddings for cross-lingual semantic search on historical Luxembourgish, a low-resource language.
-We collect historical Luxembourgish news articles spanning various time periods and use GPT-4o to segment and translate them into closely related languages, creating 20,000 parallel training sentences per language pair.
-We further create a historical bitext mining evaluation set and find that these models struggle to perform cross-lingual search on historical Luxembourgish.
-To address this, we propose a simple adaptation method using in-domain training data, achieving up to 98% accuracy in cross-lingual evaluations.
-We release our adapted models and historical Luxembourgish-German/French bitexts to support further research.
+However, pre-trained multilingual models face challenges with historical content due to OCR noise and outdated spellings.
+This study examines multilingual embeddings for cross-lingual semantic search in historical Luxembourgish (LB), a low-resource language.
+We collect historical Luxembourgish news articles from various periods and use GPT-4o for sentence segmentation and translation, generating 20,000 parallel training sentences per language pair.
+Additionally, we create a semantic search (Historical LB Bitext Mining) evaluation set and find that existing models perform poorly on cross-lingual search for historical Luxembourgish.
+Using our historical and additional modern parallel training data, we adapt several multilingual embedding models through contrastive learning or knowledge distillation and increase accuracy significantly for all models.
+We release our adapted models and historical Luxembourgish-German/French/English bitexts to support further research.
+111See https://github.com/impresso/histlux_emb for our released models, data and source code.
 Adapting Multilingual Embedding Models to Historical Luxembourgish
@@
 1 Introduction
-The ability to access and explore historical texts such as newspapers has advanced rapidly through the digitization efforts of libraries and archives.
-This exploration has traditionally been accomplished through keyword-based searches, enhanced by semantic enrichment such as named entities.
-More recently, vector-based semantic search has gained prominence through embedding models, which are also widely used in Retrieval-Augmented Generation to increase response factuality.
-In embedding benchmarks (Muennighoff et al., 2023), massively multilingual embedding models such as Multilingual E5 and GTE Multilingual (Zhang et al., 2024; Wang et al., 2024), trained on diverse multilingual corpora, have demonstrated strong performance in multilingual and cross-lingual semantic search across a wide range of languages.
-However, for low-resource languages such as Luxembourgish (LB), where multilingual models have limited exposure, their performance remains uncertain.
-When applied to semantic search in imperfectly digitized historical collections, these models must be robust to OCR errors and old spellings.
-The mismatch between this domain and the clean, contemporary texts on which these multilingual models were trained, along with their already limited support for Luxembourgish, poses a challenge for developing effective exploration tools for historical Luxembourgish newspaper archives.
-To address this issue, we compile 2,340 historical Luxembourgish news articles from different time periods and use GPT-4o to segment and translate them into contemporary German (DE) and French (FR).
+Exploration possibilities of historical texts, such as newspapers, have advanced rapidly due to digitization efforts by libraries and archives (Ehrmann et al., 2023a).
+Traditionally, tools relied on keyword-based searches, often enhanced with semantic enrichment techniques such as named entity recognition (Ehrmann et al., 2023b).
+Recent embedding benchmarks (Muennighoff et al., 2023; Enevoldsen et al., 2025) show that massively multilingual embedding models, trained on diverse multilingual corpora, perform well in both multilingual and cross-lingual semantic search.
+These models have also become integral in Retrieval-Augmented Generation (RAG), where they help retrieve more relevant and contextually appropriate documents, thereby improving the faithfulness of generated responses.
+However, for low-resource languages like Luxembourgish (LB), where multilingual models have limited exposure, their performance remains uncertain.
+Applying these models to semantic search in imperfectly digitized historical collections introduces additional challenges, as they must handle OCR errors and historical spelling variations.
+The disparity between these noisy, historical texts and the clean, modern digital-born data used to train multilingual models, combined with their limited support for Luxembourgish, complicates the development of effective exploration tools for historical Luxembourgish newspaper archives.
+To address this issue, we compile 2,338 historical Luxembourgish news articles from different time periods and use GPT-4o to segment and translate them into modern French (FR), English (EN) and German (DE).
 The resulting parallel sentences serve as fine-tuning data to adapt existing multilingual embedding models for imperfectly digitized historical Luxembourgish.
-Our contributions:
-1. To adapt multilingual embeddings for digitized historical Luxembourgish, we propose to build training data through a prompt-based translation approach with GPT-4o.
-2. For evaluation, we frame a historical Bitext Mining tasks and create a high-quality test set containing 233 source news articles (LB-DE: 2,127 sentences; LB-FR: 2,157 sentences).
-3.
-To address this task and assess our proposed adaptation method, we evaluate models based on LaBSE (Feng et al., 2022), M-GTE (Zhang et al., 2024) and LuxEmbedder (Philippy et al., 2025), and analyze potential trade-offs in performance.
-We will release our code and datasets publicly.
+Our main contributions:
+(1) We adapt multilingual embeddings for digitized historical Luxembourgish by generating training data through a prompt-based translation approach with GPT-4o.
+(2) We define a historical bitext mining task and create a high-quality cross-lingual semantic search test set with 233 source news articles (LB-DE: 2,127; LB-FR: 2,157; LB-EN: 2,105 sentences).
+(3) We fine-tune and evaluate off-the-shelf models – M-MPNet (Reimers and Gurevych, 2020), LaBSE (Feng et al., 2022), M-GTE (Zhang et al., 2024), and LuxEmbedder (Philippy et al., 2025) – to assess our adaptation methods.
+(4) We propose and evaluate a 1:1 data mixing strategy that balances noisy historical texts with clean modern texts to minimize performance degradation on modern Luxembourgish benchmarks.
 2 Related Work
+This section reviews relevant embedding models that support Luxembourgish semantic search, including monolingual Luxembourgish models and multilingual embeddings.
+Reimers and Gurevych (2020) use knowledge distillation through a strong paraphrase-trained English embedding model and parallel data to create cross-lingually aligned models.
+Multiple instances of such models have been open sourced and a particularly powerful and popular one is paraphrase-multilingual-mpnet-base-v2 (M-MPNet) which was trained on over 50 languages.
+Later within this work, we will explain how we extend this model to also support Luxembourgish.
 The multilingual bitext mining model LaBSE (Feng et al., 2022) is trained with translation ranking loss and negative samples.
@@
 One example is LuxemBERT (Lothritz et al., 2022), a monolingual BERT model pre-trained for Luxembourgish using augmented data, partially generated by translating texts from closely related languages and incorporating relevant text sources.
-Closely related to our work, LuxEmbedder (Philippy et al., 2025) used OpenAI’s text-embedding-3-small and LaBSE to bitext mine a set of 20,000 parallel sentences for each language pair between Luxembourgish, English, and French.
-These parallel sentences were then used to further fine-tune LaBSE, improving performance on contemporary Luxembourgish evaluation sets.
-However, its ability to handle Luxembourgish texts from different historical periods—potentially affected by digitization errors common in large-scale historical text collections—remains unclear.
+Closely related to our work, LuxEmbedder (Philippy et al., 2025) used OpenAI’s text-embedding-3-small and LaBSE to mine a set of parallel sentences for each pair of languages between Luxembourgish, English, and French.
+These parallel sentences (up to 20,000 per pair) were then used to further fine-tune LaBSE, improving performance on modern Luxembourgish evaluation sets.
+However, its ability to handle Luxembourgish texts from different historical periods—potentially affected by digitization errors common in large-scale historical text collections, remains unclear.
+Our work aims to extend existing embedding models to better perform cross-lingual semantic search within a collection of historical, OCR-noisy Luxembourgish texts.
+The conditions of these texts combined with the different spelling variations222Luxembourgish had no standardized spelling until 1946 and underwent multiple further reformations (eg.
+in 1999) poses an interesting generalization challenge to the models.
 3 Method
-For training and evaluating embedding models for digitized historical Luxembourgish news articles, we create parallel texts.
+To adapt and evaluate embedding models for digitized historical Luxembourgish news articles, we create parallel texts by translating them into modern German, French, and English.
+This allows the models to learn cross-lingual representations and improves their ability to align historical Luxembourgish with contemporary languages for semantic search.
 3.1 Parallel Historical Luxembourgish
-Model Historical Bitext Mining Contemporary LB Domain Adapted LBDE LBFR SIB 200 (LB) PARALUX OpenAI/text-embedding-3-small – 82.33 78.36 40.20 15.71 OpenAI/text-embedding-3-large – 88.15 86.18 58.82 26.28 LaBSE – 94.01 93.34 43.24 38.14 LuxEmbedder – 85.48 85.07 65.59 52.24 M-GTE – 87.55 85.71 55.78 70.51 LaBSE (ours) LBDE 98.52 97.98 34.02 14.70 LBFR 98.24 97.88 35.88 25.96 LBDE/FR 98.73 98.21 36.27 20.19 LuxEmbedder (ours) LBDE 98.38 97.72 50.10 25.96 LBFR 97.89 97.74 55.88 25.32 LBDE/FR 98.52 97.91 43.04 24.68 M-GTE (ours) LBDE 97.34 96.25 44.41 41.99 LBFR 96.26 95.78 59.71 50.32 LBDE/FR 97.37 96.30 49.22 44.55 Random Baseline – 0.00 0.00 14.28 50.00
-We build our translated parallel data sets LB-DE and LB-FR from monolingual Luxembourgish texts sourced from the publicly available BNL newspaper archive.111https://data.bnl.lu/data/historical-newspapers/ Our data consists of articles from newspapers published between 1841 and 1948.
-To select diverse samples for translation, we first cluster the articles into 2,000 groups by K-Mean and keep the 605 clusters with more than 20 articles.
+We build our translated parallel data sets LB-DE, LB-FR and LB-EN from monolingual Luxembourgish texts sourced from the publicly available BNL newspaper archive.333https://data.bnl.lu/data/historical-newspapers Our data consists of articles from newspapers published between 1841 and 1948.
+To select diverse samples for translation, we first cluster the articles into 2,000 groups by K-Mean on a 100-topics LDA model output444Taken from the impresso-project.ch (Ehrmann et al., 2020).
+and keep the 605 clusters with more than 20 articles.
 We select articles through a two-step process, resulting in a total of 2,340 articles, as shown in Figure 1.
-First, we retrieve the most representative article from each cluster, ensuring it contains between 5 and 20 sentences.
+First, we retrieve the most representative article from each cluster, ensuring it contains between 5 and 20 sentences (cutting of the remaining sentences).
 In a second round, we randomly sample three additional articles per cluster under the same length conditions.
-We prompt GPT-4o to segment historical Luxembourgish articles and generate sentence-level translation pairs separately for German and French (see Prompt 3).
+We prompt GPT-4o to segment historical Luxembourgish articles and generate sentence-level translation pairs separately for German, French and English (see Prompt 2).
 The model is instructed to preserve the original meaning and structure as closely as possible while reconstructing sentences affected by OCR errors that could hinder translation.
-This process yields approximately 22,500 sentence pairs for both LB-DE and LB-FR.
-Notably, the model’s consistent sentence segmentation results in 74.9% of sentences forming triplets across the three languages.
+This process yields approximately 22,500 sentence pairs for the LB-DE, LB-FR and LB-EN pair.
+Notably, GPT-4o appears to perform sentence segmentation consistently resulting in 65.0% of sentences forming exact quadruplets (4-way parallel) across the four languages.
 To ensure fidelity to the original articles, we calculate the percentage of regenerated Luxembourgish sentences that do not exactly match their source texts.
-These account for 1.5% of all historical Luxembourgish sentences per language pair, which we manually correct.
+These account for 1.4% of all historical Luxembourgish sentences per language pair, which we manually correct.
 Most mismatches result from missing or added punctuation, modernized spelling, and, in rare cases, errors caused by the LLM not adhering to the instructed format.
-To assess translation quality, a trilingual native LB speaker annotates 100 randomly sampled sentence triplets after removing 15 samples with obvious OCR issues or other factors.
-Of 100 Luxembourgish sentences (presented without context), 86 are rated comprehensible or at least confidently guessable (24).
-14 are incomprehensible in isolation due to OCR errors and old spellings.
-The translations of the incomprehensible sentences are not assessed.
-The German translations of the comprehensible and confidently guessable sentences are rated as adequate in 77 cases (89.5%): having missing details in 8 cases and only 1 case of insufficient translation.
-For the French translations, 76 cases are adequate (88.4%), 9 are missing minor details, and 1 is of insufficient quality.
-Samples from our dataset including their annotation labels are available in the Appendix (Table 2).
-3.2 Framing an Evaluation Task: Historical Bitext Mining
-From our parallel dataset, we reserve a held-out test set of 233 articles (2,127 sentences) to construct a historical LB-to-DE/FR Bitext Mining evaluation set.
-The evaluation method requires the model to assign higher similarity to the correct parallel sentence than to any of the 2,126 other candidates in the set in order to count as a true positive.
-We report bi-directionally averaged accuracy.
-To minimize false negatives stemming from near-identical sentences, we exclude candidate sentences with a Levenshtein similarity score above 0.85 to the source sentence after removing non-alphanumeric characters from both.
-This filtering affects 56 source-candidate pairs (2.6% of the source sentences) in German and 64 source-candidate pairs (2.97% of the sentences) in French.
-A human check of the excluded pairs at different thresholds confirms the appropriateness of this filtering process and the chosen threshold.
-3.3 Contemporary LB Evaluation Datasets
-To contextualize the evaluation on our targeted in-domain data, we replicate two evaluation methods of Philippy et al.
-(2025) for contemporary Luxembourgish.
-First, PARALUX is a monolingual Luxembourgish paraphrase detection test set designed to assess embedding models.
-Performance is measured by how often an embedding model assigns higher similarity to an anchor-positive pair than to an anchor-negative pair.
-The negative sentences are adversarially generated to ensure high lexical similarity, and human-corrected to ensure that they are true negatives.
-Second, SIB 200 (LB) is a repurposed dataset sourced from the ‘Flores’-dataset (Costa-jussà et al., 2022), framed as a monolingual zero-shot classification task based on embedding similarity.
-For further details, we refer to Philippy et al.
-(2025).
-3.4 Adapting Multilingual Embedding Models to Historical Luxembourgish
-We take the aforementioned LaBSE, LuxEmbedder and M-GTE and adapt them to historical Luxembourgish.
-We fine-tune these embedding models using a MultipleNegativesRankingLoss (Henderson et al., 2018) with batch size 8 over 1 epoch.
-We experiment with three fine-tuning data variants: (1) LBDE with 20,000 historical Luxembourgish sentences paired with German, (2) LBFR pairs, and (3) LBDE/FR, using 10,000 sentences from each language in mixed batches.
+To assess translation quality, a quadrilingual native speaker of Luxembourgish (LB) annotated 100 randomly selected sentence quadruplets after removing 15 samples with severe OCR problems.
+Of the 100 LB sentences presented without context, 88 were judged to be comprehensible or at least confidently guessable (23).
+The remaining 12 were considered incomprehensible due to OCR errors and archaic spellings, and their translations were not evaluated.
+For the comprehensible and confidently guessable sentences, the German translations were rated as adequate in 78 cases (88.6%), with missing minor details in 9 cases and 1 case of inadequate translation.
+The French translations showed a similar pattern: 78 were adequate (88.6%), 9 were missing minor details, and 1 was inadequate.
+The English translations were also adequate in 78 cases (88.6%), with 10 missing minor details.
+A sample of the annotated dataset is available in the appendix (Table 2).
+3.2 Framing an Evaluation Task: Historical LB Bitext Mining
+From our parallel dataset, we set aside a held-out test set of 233 articles (2,127 sentences) to establish a historical semantic search benchmark for Luxembourgish-to-German, French, and English bitext mining (LB<->DE/FR/EN).
+A prediction is considered a true positive if the embedding model assigns a higher similarity to the correct parallel sentence than to any of the 2k alternative candidates.
+We report the bidirectional average accuracy.
+To minimize false negatives caused by near-identical sentences, we exclude candidate sentences with a Levenshtein similarity score above 0.85 to the source sentence, after removing non-alphanumeric characters from both.
+This filtering affects 57 source-candidate pairs (2.7%) in German, 65 (3%) in French, and 76 (3.6%) in English.
+A human review at different thresholds confirms the appropriateness of the filtering process and the chosen threshold.
+3.3 Modern LB Evaluation Tasks
+We replicate two evaluation tasks on modern Luxembourgish from Philippy et al. (2025).
+ParaLux is a monolingual paraphrase detection test set designed to evaluate embedding models.
+Performance is measured by the proportion of cases (a total of 312 triplets) in which an embedding model assigns a higher similarity to an anchor-positive pair than to an anchor-negative pair.
+The negative sentences are adversarially generated to maintain high lexical similarity and manually verified to ensure they are true negatives.
+3.4 Adapting Multilingual Embedding Models to Historical LB
+3.4.1 Datasets
+Model
+Training Data
+Historical LB Bitext Mining
+Modern LB
+LBFR
+LBEN
+LBDE
+AVG
+SIB 200 (LB)
+ParaLux
+AVG
+Random Baseline
+00.00
+00.00
+00.00
+45.97
+14.28
+50.00
+32.14
+text-embedding-3-small
+78.36
+75.08
+82.33
+78.59
+40.20
+15.71
+27.96
+text-embedding-3-large
+86.18
+83.63
+88.15
+85.99
+58.82
+26.28
+42.55
+M-MPNet
+46.32
+45.04
+46.55
+45.97
+24.71
+26.60
+25.66
+M-MPNet(+LB Distilled)
+LBEN (Hist)
+87.23
+87.53
+89.14
+87.97
+42.65
+56.09
+49.37
+LBEN (Modern)
+75.55
+77.03
+78.09
+76.89
+59.51
+80.13
+69.82
+LBEN (Mixed)
+89.32
+89.55
+91.44
+88.79
+59.41
+80.45
+70.48
+LaBSE
+93.12
+95.27
+94.01
+94.13
+43.24
+38.14
+40.69
+LaBSE (Hist)
+LBFR
+97.73
+97.22
+98.10
+97.68
+39.61
+25.00
+32.31
+LBEN
+97.24
+97.44
+97.96
+97.54
+41.76
+22.44
+32.10
+LBDE
+97.08
+97.01
+98.52
+97.54
+34.02
+14.74
+24.38
+LaBSE (Mixed)
+LBFR
+97.40
+97.55
+98.22
+97.35
+45.69
+31.73
+47.66
+LBEN
+96.80
+97.34
+97.82
+97.75
+45.59
+36.86
+50.23
+LuxEmbedder
+84.49
+85.09
+85.48
+85.02
+65.59
+52.24
+58.92
+LuxEmbedder (Hist)
+LBFR
+97.47
+97.51
+98.24
+97.74
+50.39
+32.37
+41.38
+LBEN
+97.18
+97.29
+98.26
+97.58
+54.12
+28.85
+41.49
+LBDE
+97.25
+97.72
+98.43
+97.80
+46.76
+26.60
+36.68
+LuxEmbedder (Mixed)
+LBFR
+96.97
+97.32
+97.77
+97.72
+56.86
+38.46
+38.71
+LBEN
+97.41
+97.58
+98.26
+97.32
+56.86
+43.59
+41.23
+M-GTE
+83.68
+80.12
+87.55
+83.78
+55.78
+70.51
+63.20
+M-GTE (Hist)
+LBFR
+95.18
+94.23
+96.05
+95.15
+59.12
+57.05
+58.09
+LBEN
+95.81
+95.56
+96.52
+95.96
+54.71
+55.77
+55.24
+LBDE
+95.23
+94.61
+97.65
+95.83
+45.29
+42.31
+43.80
+M-GTE (Mixed)
+LBFR
+95.53
+95.11
+96.78
+95.80
+60.98
+60.26
+60.62
+LBEN
+95.48
+95.58
+96.55
+95.87
+67.84
+64.10
+65.70
+M-GTE (Hist, Modern: 120k)
+LBDE/FR/EN
+96.83
+97.15
+97.93
+97.30
+62.16
+62.82
+62.75
+Historical: We use 2,105 historical LB newspaper articles (excluding held-out articles) with their sentence-level translations to create a parallel training set for the following language pairs: LB-DE (20,092), LB-FR (20,010), and LB-EN (19,054) sentences.
+Modern: Philippy et al. (2025) extracted 89,405 LB-FR and 28,172 LB-EN parallel sentence pairs from RTL.lu, a trilingual news platform. This dataset was used to fine-tune the LuxEmbedder model.
+Training Data Configurations
+We investigate three data mixing strategies for model training:
+(1) Historical: 20,000 translated sentence pairs (LBFR, LBEN, or LBDE) from historical texts.
+(2) Modern: 20,000 bitext-mined sentence pairs (LBFR or LBEN) from modern Luxembourgish news.
+(3) Mixed: 20,000 Hist sentence pairs with 20,000 Modern sentence pairs in mixed batches.
+3.4.2 LB Knowledge Distillation
+We adapt M-MPNet for historical LB using multilingual knowledge distillation (Reimers and Gurevych, 2020).
+The original English model paraphrase-mpnet-base-v2 serves as a frozen teacher, while LB-EN parallel sentences are used to train the M-MPNet student model to embedded LB sentences similar to their English translations.
+We fine-tune M-MPNet for five epochs using each of the three data mixing strategies: (1) Historical, (2) Modern, and (3) Mixed.
+3.4.3 Contrastive Loss
+We adapt LaBSE, LuxEmbedder, and M-GTE to historical LB using contrastive learning.
+Specifically, we fine-tune these embedding models using MultipleNegativesRankingLoss (Henderson et al., 2017), with a batch size of 8 for one epoch.
+For fine-tuning, we apply two of the previously defined data mixing strategies: (1) Historical and (3) Mixed.
 4 Results
-We show the results of various models in Table 1.
-Historical Bitext Mining – Cross-Lingual Semantic Search
-Among off-the-shelf models, LaBSE is the strongest model in both languages.
-Surprisingly, LuxEmbedder, the Luxembourgish-tuned version of LaBSE, shows an 8.5pp performance drop for both language pairs in our bitext mining task, despite improved performance on contemporary Luxembourgish tasks.
-Similarly, M-GTE underperforms LaBSE by 5.5pp.
-Both OpenAI embedding models demonstrate moderate performance in all tasks.
-For domain-adapted models, performance improves significantly in historical bitext mining for both language pairs after training on 20,000 parallel sentences, whether using LBDE or LBFR data.
-However, the best results are consistently achieved with the 20,000 LBDE/FR sentence pairs, yielding higher performance across all models.
-Notably, after domain adaptation, LuxEmbeddder nearly matches the adapted LaBSE, achieving over 98% accuracy and closing the performance gap observed in the off-the-shelf models.
-In contrast, M-GTE, the general embedding model, appears to plateau at lower performance levels, trailing the specialized domain-adapted bitext mining model by 1.5pp.
-Contemporary vs. historic Luxembourgish: A notable trade-off
-While our domain-adapted models improve the embedding of historical Luxembourgish for cross-lingual semantic search in contemporary German and French (as shown on the right side of the Table 1), their performance on monolingual contemporary tasks, particularly PARALUX, declines noticeably after domain adaptation.
-Interestingly, this degradation is more pronounced when models are trained on the linguistically closer LBDE pairs than LBFR pairs.
-The learning curves in Figure 2 illustrate the continuous decline in monolingual performance.
-Notably, most cross-lingual performance gains occur within the first 2,500 samples, before monolingual Luxembourgish performance declines more sharply.
-This highlights the trade-off between optimizing for historical cross-lingual tasks and maintaining monolingual performance, as seen in the weaker monolingual results of the best off-the-shelf LaBSE model.
+Table 1 shows the performance of the off-the-shelf and adapted models on the historical Luxembourgish bitext mining and the modern LB evaluation tasks.
+Among the off-the-shelf(in cursive) models, LaBSE is the strongest model in all three languages.
+Surprisingly, LuxEmbedder, a LB-tuned version of LaBSE, shows an average performance drop of 9pp across language pairs in our bitext mining task, despite improved performance on the modern LB tasks.
+Similarly, M-GTE underperforms LaBSE by 10.4pp.
+Both OpenAI embedding models (text-embedding-3-small/large) show moderate performance.
+Among the models contrastively adapted using the Hist pairs, the performance in historical bitext mining improves significantly, reaching up to 97.8% accuracy.
+Notably, after domain adaptation, LuxEmbedder matches the adapted LaBSE, reaching over 97.8% accuracy and closing the performance gap observed in the standard models.
+Meanwhile, the customized M-GTE models lag behind by about 2pp.
+Interestingly, across all model architectures, training on any language pair improves performance similarly across all pairs, showing a positive cross-lingual transfer.
+These models experience significant performance drops in modern LB evaluations, particularly in ParaLux.
+However, adapting these models with mixed batches of Hist and Modern sentence pairs partially mitigated performance loss on the Modern LB evaluation tasks.
+Within LaBSE and M-GTE, this adaptation even improved the performance on SIB-200 topic classification, while sacrificing only up to 1% of the performance on historical bitext mining.
+These mixed-data adapted models provide an overall stronger general backbone for cross-lingual semantic searching within collections containing both historical and modern Luxembourgish.
+The M-MPNet model, before distilling EN-LB knowledge, performs poorly in all LB evaluations, despite its proven exact matching capabilities in other languages, confirming its lack of support for the language.
+After distilling LB with any dataset, the model performs magnitudes better across the board.
+When distilled with a single dataset, the model performs best on historical semantic search tasks with the Hist sentence pairs.
+In contrast, when distilled with Modern sentence pairs, the model excels on modern LB tasks, achieving 80% accuracy on ParaLux555As shown in recent work (Michail et al., 2025), results on adversarial paraphrase discrimination test sets might not accurately reflect performance on semantic search in general.
+Therefore, this result should be interpreted with caution.
+and outperforming the second-best M-GTE, which achieves 70%.
+Finally, distilling with the mixed data set yields the best results in all evaluations, demonstrating the synergy between the two sources.
+However, even the Mixed-data distilled M-MPNet model only achieves an average accuracy across pairs of 90% in historical bitext mining, trailing the contrastive domain-adapted LaBSE and LuxEmbedder by 8pp and the off-the-shelf LaBSE model by 4pp.
+The Final Model: Mix it All
+For a final all-purpose model covering both historical and modern LB, we contrastively adapt M-GTE to all language pairs of Hist while preserving an equal number of Modern sentence pairs, regardless of language.
+The adaptation dataset consists of 20,000 LB-FR/EN/DE (Hist), 20,000 LB-EN (Modern), and 40,000 LB-FR (Modern), for a total of 120,000 sentence pairs.
+It is the best-performing historical semantic search M-GTE model, achieving an average accuracy of 97.5% across all language pairs.
+This model outperforms the adapted LaBSE and LuxEmbedder models on SIB-200 (+6pp) and ParaLux (+20pp), while performing similarly to them in the historical bitext mining evaluations.
 5 Conclusions
-In this work, we present a case study on embedding imperfectly digitized historical Luxembourgish texts.
-As off-the-shelf models prove suboptimal for this task, we propose a method to synthesize reliable data for adapting embedding models to this domain.
-We generate parallel sentence-segmented documents by translating historical Luxembourgish newspaper articles into two related higher-resource languages (German and French) using GPT-4o as a machine translation system.
-To establish a hold-out test set, we reserve 233 articles and frame a bitext mining task for retrieving historical Luxembourgish sentences.
-Evaluating various multilingual embedding models on this task, we demonstrate that our adaptation approach improves performance by up to 13 percentage points.
-While our adapted models perform effectively across different historical periods, this improvement comes at the cost of reduced performance in natural language understanding (NLU) for clean contemporary texts.
-This trade-off highlights the need for future research into embedding historical digitized texts affected by significant OCR errors.
+In this work, we explore the adaptation of multilingual embedding models to digitized historical LB texts, a task where off-the-shelf models struggle due to limited exposure and their reliance on clean modern data.
+To address this issue, we generate parallel sentence-segmented documents by translating historical Luxembourgish newspaper articles into French, English, and German using GPT-4o.
+To evaluate the effectiveness of adaptation, we design a historical bitext mining task with a held-out test set of 233 articles.
+Our results show that adaptation to parallel historical data improves retrieval accuracy by up to 13pp.
+However, this adaptation introduces trade-offs, particularly reducing performance on modern LB tasks that require high semantic precision, such as adversarial paraphrase detection.
+We mitigate this problem through a balanced data mixing strategy that helps preserve modern LB performance while improving historical text semantic search capabilities.
+These results demonstrate the effectiveness of domain adaptation for historical text processing and suggest that such approaches could benefit low-resource languages facing digitization challenges.
+Such improvements are particularly relevant for libraries and archives, where effective cross-lingual semantic search can improve the discoverability of historical documents and support digital exploration.
 Limitations
-Our findings present both positive and negative aspects.
+Our findings paves the way for better semantic search systems within Luxembourgish archives.
 On the one hand, our method demonstrates clear benefits for the targeted use case, effectively embedding heterogeneous digitized historical texts and revealing shortcomings in off-the-shelf models.
-On the other hand, performance drops in control tasks on contemporary texts raise concerns.
-Does improved robustness to digitization errors come at the cost of weaker language understanding and reduced sensitivity to semantic nuances?
-Can future training strategies mitigate this trade-off?
-These important questions remain open for future research.
-Acknowledgements
+Through our exploration of adaptation methodologies, we have produced practical embeddings for semantic search while mostly preserving modern LB performance.
+On the other hand, we have not strictly reached a single best model across all evaluation sets.
+For example, in all of our adapted models, performance on ParaLux declines, possibly indicating interference with modern LB understanding and reduced sensitivity to semantic nuances.
+Overall, we have applied a single adaptation method for each model type across all available data mixes, ensuring alignment with the models’ initial training methods.
+Exploring alternative adaptation approaches may reveal additional patterns.
+One problem with our evaluation is that they are all at the sentence level, whereas applications of such models would often be at the paragraph, article, or document level.
+The hypothesis that our improved performance would be reflected when embedding longer segments of text is possible, but not guaranteed.
+Lastly, while our research focuses on historical Luxembourgish, our methodology may also be useful for developing semantic search models in other underrepresented languages, which we do not examine in this study.
+Acknowledgments
+We would like to thank Fred Phillipy for helping with the human annotation.
 This research is conducted under the project Impresso – Media Monitoring of the Past II Beyond Borders: Connecting Historical Newspapers and Radio.
@@
 Appendix A Appendix
-| Newspaper | Year | Sentence | LB Comprehensibility Rating | DE Translation Rating | FR Translation Rating |
+| Newspaper | Year | Sentence | LB Compr. | FR Compr. | EN Compr. | DE Compr. |
 | D’Union | 1946 |
 LB: Si bestét aus 18 000 tuben, weit 30 tonnen a kascht 400 000 dollar.
-DE: Sie besteht aus 18.000 Röhren, wiegt 30 Tonnen und kostet 400.000 Dollar. FR: Elle est composée de 18 000 tubes, pèse 30 tonnes et coûte 400 000 dollars. |
-Comprehensible | Adequate | Adequate |
+FR: Elle est composée de 18 000 tubes, pèse 30 tonnes et coûte 400 000 dollars.
+EN: It consists of 18,000 tubes, weighs 30 tons, and costs 400,000 dollars.
+DE: Sie besteht aus 18.000 Röhren, wiegt 30 Tonnen und kostet 400.000 Dollar.
+|
+Comprehensible | Adequate | Adequate | Adequate |
 | Ons Jorgen | 1946 |
 LB: „Daf net, mais ech hu kc properen Teller me’,"
-DE: „Das nicht, aber ich habe keinen sauberen Teller mehr. FR: "Certainement pas, mais je n’ai plus d’assiette propre," |
-Comprehensible | Adequate | Adequate |
+FR: "Certainement pas, mais je n’ai plus d’assiette propre," EN: "Not really, but I don’t have a single clean plate anymore." DE: „Das nicht, aber ich habe keinen sauberen Teller mehr. |
+Comprehensible | Adequate | Adequate | Adequate |
 | De Letzeburger | 1893 |
 LB: De Batti: Elo hätte mer d’Stemmung gut eriwer, hätte mer elo och nach Rén.
+FR: Le Batti : Maintenant, nous aurions bien passé l’ambiance, si seulement nous avions aussi encore Rén.
+EN: Batti: Now we would have a good atmosphere if we also had some rain.
 DE: Der Batti: Jetzt hätten wir die Stimmung gut geschafft, hätten wir jetzt auch noch Regen.
-FR: Le Batti : Maintenant, nous aurions bien passé l’ambiance, si seulement nous avions aussi encore Rén.
 |
-Confidently Guessable | Adequate | Adequate |
+Confidently Guessable | Adequate | Adequate | Adequate |
 | De Letzeburger | 1905 |
 LB: Op d’Weis: Das ist im Lehen hfisslich eingerichtet.
-DE: Zur Melodie: Das ist im Leben hässlich eingerichtet. FR: À la manière de : Cela est arrangé vilaine dans la vie. |
-Confidently Guessable | Adequate | Adequate |
+FR: À la manière de : Cela est arrangé vilaine dans la vie. EN: To the tune: It is poorly arranged in life. DE: Zur Melodie: Das ist im Leben hässlich eingerichtet. |
+Confidently Guessable | Adequate | Adequate | Adequate |
 | De Letzeburger | 1893 |
 LB: Wann och d’Liss’ché wéss ze feischtren.
-DE: Wenn auch die Liss’ché weiß zu feilschen. FR: Même si Liss’ché sait lutter. |
-Incomprehensible | / | / |
+FR: Même si Liss’ché sait lutter. EN: Even if Lisette knows how to flirt. DE: Wenn auch die Liss’ché weiß zu feilschen. |
+Incomprehensible | / | / | / |
 | Obermosel-Zeitung | 1924 |
 LB: In vielen vorkern Bincl alle Krank, 80 cla,BB «lie ?eläer nicdt deBtellt terrien Können.
-DE: In vielen Dörfern sind alle krank, so dass die Felder nicht bestellt werden können. FR: Dans de nombreux villages, tous sont malades, si bien que les champs ne peuvent pas être cultivés. |
-Incomprehensible | / | / |
+FR: Dans de nombreux villages, tous sont malades, si bien que les champs ne peuvent pas être cultivés.
+EN: In many places, all are sick, so that the fields cannot be tended.
+DE: In vielen Dörfern sind alle krank, so dass die Felder nicht bestellt werden können.
+|
+Incomprehensible | / | / | / |